Grade 12Physical Chemistry

How many grams of NaBr must be added to 270g of water to lower the vapour pressure by 3.125mmHg.at a temperature at which vapour pressure of water is 50mmHg (at. Mass of Na =23 , Br= 80).Assume 100% ionisation of NaBr
a.151.5g. b.0.51g. C. 5.5g. D.1.5g. E.51.5g

Bhavya
7 Years agoGrade 12

1 Answer

Samyak Jain
7 Years ago
According Roult’s law, for non-electrolyte solutions
\DeltaP / P0  =  XB, where XB is the mole fraction of solute in solution = nB / (nA + nB)
A represents solvent and B represents solute.
For dilute solutions, XB \approx nB / nA .
For dilute electrolyte solutions, the formula becomes
\DeltaP / P0  =  i.nB / nA , where i is Vant Hoff factor.
Here i = 2, P0 = 50 mm Hg, \DeltaP = 3.125 mm Hg, nA = 270 g / (18 g/mol) = 16 mol.
\therefore 3.125 / 50  =  2.nB / 16  or  0.0625 = nB / 8
So, nB  =  0.5  =  Required mass of NaBr / Molar mass of NaBr
Thus, Required mass of NaBr = 0.5 x 103 = 51.5 g
Option E is correct.